OpenStudy (anonymous):

Simplify 2 over quantity x squared plus 8 x plus 15 end quantity plus 1 over quantity x squared plus 11 x plus 30 (1 point)

OpenStudy (phi):

I would try to factor each of the quadratics as a first step to finding a common denominator

OpenStudy (anonymous):

will you try it and tell me what to get because i can't seem to figure out anything

jhonyy9 (jhonyy9):

2 1 ------------- = --------------- x^2 +8x +15 x^2 +11x +30 so you know that 8= 5+3 15 = 5*3 and 11 = 6+5 30 = 6*5 can you make it now ?

OpenStudy (anonymous):

so after i factor it what do you do

jhonyy9 (jhonyy9):

read what have wrote in the first line above @phi ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

actually not really. i really hate math and have no idea what to do. i factored it thats it

OpenStudy (phi):

what do you have so far ?

OpenStudy (anonymous):

((2)/(x+5)(x+3))+((1)/(x+6)(x+5))

OpenStudy (phi):

next you need a common denominator. both have (x+5) I would use (x+5)(x+3)(x+6) to get that denominator for the first fraction, multiply top and bottom by (x+6) (but don't multiply it out, just write it down)

OpenStudy (phi):

for the second fraction, multiply top and bottom by (x+3)

OpenStudy (anonymous):

ok now what

OpenStudy (phi):

what do you have so far ?

OpenStudy (anonymous):

((2)(x+6)/(x+5)(x+3)(x+6))+((1)(x+3)/(x+6)(x+5)(x+3))

OpenStudy (phi):

now both fractions have the same denominator, so that will be the denominator of the answer. concentrate on just the tops: 2(x+6) + 1(x+3) I would distribute the 2 in the first part, and simplify. can you do that ?

OpenStudy (anonymous):

2x+12+x+3=2x^2+15

OpenStudy (phi):

when you add, you can't get x^2 12+3 is 15 so that is ok 2x + x or (2x + 1x) means : you have 2 x's and add 1 more x. how many x's do you have?

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

3x+15

OpenStudy (phi):

so the top 2(x+6) + 1(x+3) becomes 3x+15 can you factor 3 out of this ?

OpenStudy (anonymous):

x+5

OpenStudy (phi):

you write 3x+15 as 3(x+5) so now you have the top as 3(x+5) if we put that over the bottom, what do we get ?

OpenStudy (anonymous):

3(x+5)/((x+6)(x+5)(x+3))

OpenStudy (phi):

notice you have (x+5) in the top and bottom. anything divided by itself is 1, so you can simplify your answer.

OpenStudy (anonymous):

3/(x+6)(x+3))

OpenStudy (phi):

Finally, because anything divided by itself is 1 *except 0*, we do should make a note that x ≠ -5 because in the original, that would give us 0/0 which is undefined. the final answer is \[ \frac{3}{(x+6)(x+3)} \text{ , x≠ -5 } \] x cannot be -6 or -3 either (no divide by 0), but we can see that just looking at the fraction

OpenStudy (anonymous):

ok thanks so much for your help